A travel agency offers two Boston outings. Plan A includes hotel accommodations for three nights and two pairs of baseball tickets worth $556.40. Plan B includes hotel accommodations for five nights and four pairs of baseball tickets worth $973.
Let x represent the cost of one night's hotel accommodation and let y represent the cost of one pair of baseball tickets. Write a system of equations you could solve to find the cost of one night's hotel accommodation and one pair of baseball tickets.
Solve the system.
The yellow pages identify two different local electrical businesses. Business A charges $50 for a service call, plus an additional $39 per hour for labor. Business B charges $32 for a service call, plus an additional $45 per hour for labor.
Let x represent the number of hours of labor and let y represent the total charge. Write a system of equations you could solve to find the length of a service call for which both businesses charge the same amount.
Solve the system.
Which company would you use? Why?

Answer by mananth(16949)

A travel agency offers two Boston outings. Plan A includes hotel accommodations for three nights and two pairs of baseball tickets worth $556.40. Plan B includes hotel accommodations for five nights and four pairs of baseball tickets worth $973.
Let x represent the cost of one night's hotel accommodation and let y represent the cost of one pair of baseball tickets. Write a system of equations you could solve to find the cost of one night's hotel accommodation and one pair of baseball tickets.
Solve the system.
3x+2y=556.40......1
5x+4y=973..........2
multiply (1) by -2
-6x-4y=-1112.80
5x+4y=973
add them up
-x=-139.8
x= 139.8 one night stay.
plug value of y in (2)
139.8*5+4y=973
y = 68.50 price of ticket


The yellow pages identify two different local electrical businesses. Business A charges $50 for a service call, plus an additional $39 per hour for labor. Business B charges $32 for a service call, plus an additional $45 per hour for labor.
Let x represent the number of hours of labor and let y represent the total charge. Write a system of equations you could solve to find the length of a service call for which both businesses charge the same amount.
Solve the system.
Which company would you use? Why?
50+39x=32+45x
6x=18
x=3 hours